The xyz manufacturing bulbs hypothesis that the life of bulbs is 145 hours with a known standard deviation of 210 hours. A random sample of 25 bulbs give A mean life of 130 hours. Using 0.05 level of significance, can the co. Conclude that the mean life of bulbs is less than 145 hours?
1
Expert's answer
2020-04-03T15:05:24-0400

Null hypothesis "H_0 :\mu=145."

Alternative hypothesis "H_a: \mu<145".

Test statistic: "z=\frac{\bar x-\mu}{\frac{\sigma}{\sqrt{n}}}=\frac{130-145}{\frac{21}{\sqrt{25}}}=-3.57."

P-value: "p=0.0002."

Since the P-value is less than 0.05, reject the null hypothesis.

We can conclude that the mean life of bulbs is less than 145 hours.